OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (anonymous):

@zepdrix Hey!! I remember you helped me with a question similar to this before!I get the basic idea of it but I forgot how to turn the s(t) setup into the normal like function, like -6t is after -2? can you help me out with how I set this up to solve :D

zepdrix (zepdrix):

Hmm I forget, Have you learned the shortcuts for taking derivatives? Or do we need to use the `Limit Definition of the Derivative`?

OpenStudy (anonymous):

Usually I was using limh->0 f(x+h)-f(x)/h :D

OpenStudy (anonymous):

ds/dt = -6

zepdrix (zepdrix):

\[\Large \lim_{h\to0}\frac{s(t+h)-s(t)}{h}\]Using limit? Ok cool. \[\Large s(t)=-2-6t\]\[\Large s(t+h)=-2-6(t+h)\]

zepdrix (zepdrix):

Oh oh right, they said at t=2.

OpenStudy (anonymous):

the t value doesn't matter, because the object is moving at constant velocity

zepdrix (zepdrix):

So we actually wanna use, \[\Large s(2)=-2-6(2)\]\[\Large s(2+h)=-2-6(2+h)\] \[\Large s'(2)=\lim_{h\to0}\frac{s(2+h)-s(2)}{h}\] Understand how to plug these in and simplify? :o

OpenStudy (anonymous):

umm not really :( I never used it that way before with 2 in () afte s

zepdrix (zepdrix):

Function notation is tricky D: Gotta try and get comfortable with it. So like...

zepdrix (zepdrix):

Limit definition of the derivative: \[\Large f'(x)=\lim_{h\to0}\frac{f(x+h)-f(x)}{h}\] But we can evaluate it at specific points.\[\Large f'(3)=\lim_{h\to0}\frac{f(3+h)-f(3)}{h}\] Is that the part that's confusing you? :o

OpenStudy (anonymous):

no when it's f(x) I completely understand but now that it's s(t) and the -6t comes after the -2 I'm in a brain tizzy :'(

zepdrix (zepdrix):

We don't have to call it s(t) if that's confusing. We can instead call our function,\[\Large s(t)=-2-6t \qquad\to\qquad f(x)=-2-6x\]

zepdrix (zepdrix):

Let's try to fill in the pieces.\[\Large \color{royalblue}{f(x)=-2-6x}\]\[\Large \color{orangered}{f(x+h)=-2-6(x+h)}\] \[\large f'(x)=\lim_{h\to0}\frac{\color{orangered}{f(x+h)}-\color{royalblue}{f(x)}}{h} \quad\to\quad \lim_{h\to0}\frac{\color{orangered}{-2-6(x+h)}-\color{royalblue}{(-2-6x)}}{h}\]

OpenStudy (anonymous):

ahhh so -2-6x-6h+2+6x/h !?

zepdrix (zepdrix):

mm yah that sounds right.

zepdrix (zepdrix):

Keep simplifying, it should clean up nicely.

OpenStudy (anonymous):

so then we have -6h/h and h and h cancel out so the answer is -6?

zepdrix (zepdrix):

\[\Large f'(x)=-6 \qquad\to\qquad s'(t)=-6\]Ok good, but they wanted you to evaluate the function at t=2, what do you get? ;)

OpenStudy (anonymous):

-3!??...(:?

zepdrix (zepdrix):

\[\Large s'(2)=?\]Plug 2 into all of your t's in the equation :X *giggles*

OpenStudy (anonymous):

wait so am I still keepin -6?? oh lawwwwd :) is it -12?

OpenStudy (anonymous):

OH into the original!?

zepdrix (zepdrix):

There are no t's, right? :o The function is just a constant -6 at all times. \[\Large s'(2)=-6\]

OpenStudy (anonymous):

YOU TRICKED ME!!! xD LOL that's why I was confused xDDD

OpenStudy (anonymous):

you're evil :3

zepdrix (zepdrix):

hah :3

OpenStudy (anonymous):

but thank yuhh <3 @zepdrix :D